# The Role of the Schwinger Effect in Superradiant Axion Lasers

### v2

@@ Appendix A Behavior of functions of $\chi$
$\chi-\hat{\rho}_{0}=\sqrt{{\hat{T}_{e^{\pm}}}^{2}-1}-\tan^{-1}\sqrt{{\hat{T}_{e^{\pm}}}^{2}-1},$ (53)
@@
This makes it clear that $\frac{\Gamma_{e^{\pm}}}{\Gamma_{\gamma}}$ is a function only of $\chi$. Its asymptotic behavior may be derived by noting that, in the limit $\chi\to\infty$, the arctangent term in eq. (53) becomes insignificant, and so
$\Gamma_{e^{\pm}}\xrightarrow[\chi\to\infty]{}\frac{\chi^{3}}{6\int_{0}^{\chi}\sqrt{\left(\chi-\hat{\rho}_{0}\right)^{2}+1}\hat{\rho}_{0}\textrm{d}\hat{\rho}_{0}}\Gamma_{\gamma},$ (55)
which may be evaluated to find that $\Gamma_{e^{\pm}}\to\Gamma_{\gamma}$, which is also what one would expect physically.

### v3

@@ Appendix B Behavior of functions of $\chi$
$\chi-\hat{\rho}_{0}=\sqrt{\hat{T}_{e^{\pm}}\left(2+\hat{T}_{e^{\pm}}\right)}+\frac{\pi}{3}-2\tan^{-1}\frac{\sqrt{3}\left(1+\hat{T}_{e^{\pm}}\right)+2\sqrt{\hat{T}_{e^{\pm}}\left(2+\hat{T}_{e^{\pm}}\right)}}{3+\hat{T}_{e^{\pm}}},$ (132)
@@
This makes it clear that $\frac{\Gamma_{e^{\pm}}}{\Gamma_{\gamma}}$ is a function only of $\chi$. Its asymptotic behavior may be derived by noting that, in the limit $\chi\to\infty$, the arctangent term in eq. 132 becomes insignificant, and therefore
$\Gamma_{e^{\pm}}\xrightarrow[\chi\to\infty]{}\frac{\chi^{3}}{6\int_{0}^{\chi}\left(\sqrt{\left(\chi-\hat{\rho}_{0}\right)^{2}+1}-1\right)\hat{\rho}_{0}\textrm{d}\hat{\rho}_{0}}\Gamma_{\gamma}.$ (134)
This may be evaluated to find that
$\lim_{\chi\to\infty}\Gamma_{e^{\pm}}=\Gamma_{\gamma},$ (135)
which is what one would expect physically.